a car rental company charges $21 per day and 10 cents per mile. if a car was rented for 2 days, how far was it driven if the total came to $220.00?
Answer by checkley77(12844)

2*21+.10x=220
42+.10x=220
.10x=220-42
.10x=178
x=178/.10
x=1,780 miles were driven.
proof:
2*21+.1*1,780=220
42+178=220
220=220